OpenStudy (anonymous):

(2x)^-6 Simplify. Use positive exponents only.

OpenStudy (anonymous):

What does it mean by the "use positive exponents only"? When I get my answer it looks like 1/64(x^-6)

OpenStudy (anonymous):

1/64x^6

OpenStudy (anonymous):

How come it isn't -6?

OpenStudy (anonymous):

you multiplied wrong i think.

jimthompson5910 (jim_thompson5910):

\[\large (2x)^{-6} = \frac{1}{(2x)^6}\] \[\large (2x)^{-6} = \frac{1}{2^6*x^6}\] \[\large (2x)^{-6} = \frac{1}{64x^6}\]

jimthompson5910 (jim_thompson5910):

I'm using the idea that \[\large x^{-k} = \frac{1}{x^k}\]

OpenStudy (anonymous):

Okay that makes more sense. Thank you!

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

would it be 1/64(x^6)

jimthompson5910 (jim_thompson5910):

It is \[\large \frac{1}{64x^6}\]

OpenStudy (anonymous):

this might be a stupid question but cant the exponent be distributed later on to the 64?

jimthompson5910 (jim_thompson5910):

it already has "distributed" that's how (2x)^6 turned into 64x^6

jimthompson5910 (jim_thompson5910):

when you have 64x^6, the exponent of 6 only applies to the x

OpenStudy (anonymous):

gotcha, thank ya

jimthompson5910 (jim_thompson5910):

yw